# Mesh2SLAM in VR: A Fast Geometry-Based SLAM Framework for Rapid Prototyping in Virtual Reality Applications

### v1

@@ Introduction
Abstract - Simultaneous Localization and Mapping (SLAM) is a foundational method with broad applications in robotics and computer vision, critical for the operation of augmented and virtual reality (AR/VR) devices. SLAM algorithms rely on various sensors, such as LIDAR, cameras, and depth sensors, to build a map of the environment while simultaneously estimating the device’s location. Visual SLAM (V-SLAM) [^1] This work focuses on V-SLAM specifically, for simplicity, it will be referred to as SLAM throughout this document. primarily uses images as input data. However, additional processing of the image stream is required to detect and match salient features between frames [^2] Considering indirect, also known as feature-based V-SLAM methods.. Accurate detection and matching of these features are vital for estimating camera poses between consecutive frames, known as localization.
@@
  • First, our approach eliminates feature association errors, enhances position estimation accuracy and boosts runtime efficiency, making it also a potential candidate for generating ground-truth data in related applications.
@@ 2.2 Frame Capture
Complementarily, z components, representing depth with respect to the camera, are kept in view space (2) in order to assert depth thresholds. In this case, depth can be used to include only vertexes that are within acceptable range.
The values for clip and view can now be used to compute the normalized screen coordinates (ndc):
$x_{\text{ndc}}$ $=$ $\mathbf{v}_{\text{clip}}.x\cdot\mathbf{v}_{\text{clip}}.w^{-1}$
$y_{\text{ndc}}$ $=$ $\mathbf{v}_{\text{clip}}.y\cdot\mathbf{v}_{\text{clip}}.w^{-1}$
$z$ $=$ $-\mathbf{v}_{\text{view}}.z$
And the screen coordinates u, v:
@@
Together with the vertex’s unique identifier i, a frame captured contains all features visible within camera view frustum and maximum depth threshold:
@@ 3.3 Virtual Reality Practical Evaluation
4 Limitations
Mesh2SLAM is limited to virtual environments, which restricts its applicability to real-world scenarios. Additionally, unbalanced spatial density or clustering of mesh vertices can impact tracking performance and pose estimation accuracy.
While our evaluation was conducted using the Meta Quest 2, further testing across a wider range of HMD specifications would likely yield similar results. However, such testing would contribute to a more comprehensive evaluation.
@@ 5 Conclusion
In this work, we presented Mesh2SLAM, a fast and efficient geometry-based SLAM framework for fast prototyping that operates on HMDs for virtual reality. We have demonstrated its main novelty by directly utilizing mesh geometry components, vertex features, and leveraging GPU acceleration, thereby enhancing position estimation accuracy and drastically reducing computational overhead. To our knowledge, it is the first real-time SLAM method running on virtual environment as a standalone application, without tethering, on low-budget off-the-shelf HMDs.
We believe, based on the described evaluations and demonstrated potential, that Mesh2SLAM offers a novel and unconventional approach to SLAM. Its high performance and accuracy make it ideal for prototyping and also to generate ground-truth data in virtual scenarios. By simplifying SLAM and reducing computational demands, it enables researchers to efficiently test and refine concepts, advancing AR/VR research.

### v4

@@ Introduction
Simultaneous Localization and Mapping (SLAM) is a foundational method with broad applications in robotics and computer vision, critical for the operation of augmented and virtual reality (AR/VR) devices. SLAM algorithms rely on various sensors, such as LIDAR, cameras, and depth sensors, to build a map of the environment while simultaneously estimating the device’s location. Visual SLAM (V-SLAM) [^1] This work focuses on V-SLAM specifically, for simplicity, it will be referred to as SLAM throughout this document. primarily uses images as input data. However, additional processing of the image stream is required to detect and match salient features between frames [^2] Considering indirect, also known as feature-based V-SLAM methods.. Accurate detection and matching of these features are vital for estimating camera poses between consecutive frames, known as localization.
@@
  • First, our approach eliminates feature association errors, enhances position estimation accuracy and boosts runtime efficiency, making it also a potential candidate for use as ground-truth in related applications.
@@ 2.2 Frame Capture
Complementarily, z components, representing depth with respect to the camera, are kept in view space (2) in order to assert depth thresholds; Depth can be used to include only vertexes that are within acceptable range.
The values for clip and view can now be used to compute the normalized screen coordinates (ndc) (3a),(3b) and depth (3c):
$x_{\text{ndc}}$ $=\mathbf{v}_{\text{clip}}.x\cdot\mathbf{v}_{\text{clip}}.w^{-1},$ (3a)
$y_{\text{ndc}}$ $=\mathbf{v}_{\text{clip}}.y\cdot\mathbf{v}_{\text{clip}}.w^{-1},$ (3b)
$z$ $=-\mathbf{v}_{\text{view}}.z.$ (3c)
And finally screen coordinates u, v are obtained with w and h as width and height of screen resolution:
@@
As a result, a vertex feature aggregates the screen coordinates along with the descriptor, the vertex’s unique identifier i:
@@ 3.3 Virtual Reality Practical Evaluation
4 Scope and Constraints
Mesh2SLAM is limited to virtual environments, which restricts its applicability to real-world scenarios. Additionally, unbalanced spatial density or clustering of mesh vertices can impact tracking performance. Although the scene is currently treated as a single mesh, future extensions could handle large-scale environments with multiple meshes, which may be achieved through other optimization, sorting, or culling mechanisms. While our evaluation focused on the Meta Quest 2, we expect testing on other HMDs will likely yield similar results and provide a more comprehensive assessment.
@@ 5 Conclusion
In this work, we presented Mesh2SLAM, a fast and efficient geometry-based SLAM framework for prototyping that operates on HMDs in virtual reality. Its main novelty lies in utilizing mesh geometry components, vertex features, and leveraging GPU acceleration to enhance position estimation accuracy and drastically reduce computational overhead. To our knowledge, it is the first real-time SLAM method running in a virtual environment as a standalone application, without tethering, on low-budget, off-the-shelf HMDs.
As a SLAM simulation, Mesh2SLAM maintains SLAM functionality while bypassing issues in image-based methods that hinder prototyping. In the context of XR development, its lightweight sparse map representation is well-suited for localization-focused tasks, including tracking user movement, supporting spatial interactions, validating pose estimation methods, and benchmarking other localization algorithms. Additionally, future considerations include extending it to collaborative localization or multi-agent SLAM. These capabilities make it a valuable tool for efficiently testing and refining concepts, driving advancements in XR research.